# SpikeStereoNet: A Brain-Inspired Framework for Stereo Depth Estimation from Spike Streams

### v1

@@ 1 Introduction
Depth perception is fundamental for navigating and interacting with the 3D world, driving applications from robotic manipulation [1, 2, 3, 4] to autonomous navigation [5, 6, 7, 8, 9, 10, 11]. Traditional stereo vision estimates depth from calibrated image pairs captured by frame-based cameras; however, it suffers from motion blur and latency in dynamic scenes. Biological vision systems efficiently process visual inputs using sparse, asynchronous spikes, achieving remarkable speed and energy efficiency [12, 13, 14, 15, 16, 17, 18]. Neuromorphic spike cameras [19, 20, 21, 22, 23, 24, 25, 26, 27] implement this principle in hardware, delivering ultra-high temporal resolution (up to 40,000 Hz) and capturing rich luminance information through asynchronous binary spike streams. This makes them suitable for robust perception tasks, such as stereo depth estimation, particularly in highly dynamic scenarios where conventional methods are ineffective.
@@
Despite their potential, spike cameras present distinct challenges for stereo depth estimation. Their asynchronous, binary, high-throughput streams conflict with frame-based algorithms that expect synchronous, intensity-valued image pairs [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42], and any conversions to frames introduce temporal quantization errors, motion blur, or significant computational overhead [23, 25, 43], diminishing the sensor’s intrinsic advantages. Although event-based stereo methods have been developed for DVS cameras (e.g. [44, 45, 46, 47, 48]), they rely on temporal contrast, whereas spike cameras emit integrated intensity streams that require tailored processing techniques. Critically, the field lacks specialized algorithms and benchmarks designed for stereo depth estimation directly from raw spike streams, hindering progress in this promising field.
@@
To bridge this gap, we introduce SpikeStereoNet (Fig. 1), an end-to-end, biologically inspired framework for generating stereo depth directly from raw spike streams. SpikeStereoNet integrates a recurrent spiking neural network (RSNN) into an iterative refinement loop [49] and models neuronal interactions [50] to capture the spatiotemporal dynamics inherent in spike data. The analysis of neuronal dynamics confirms the temporal stability, convergence, and expressive feature separation of the RSNN components. We introduce two novel benchmark datasets: a large-scale synthetic dataset with diverse scenes and ground-truth depth, and a real-world dataset containing synchronized stereo spike streams and corresponding depth sensor measurements. Experimental results demonstrate that SpikeStereoNet generalizes effectively, maintaining robust performance even when trained with substantially limited data. Our main contributions are as follows:
@@ 2 Related Work
Spike Cameras and its Applications. Neuromorphic cameras, including event cameras [51, 52, 53] and spike cameras [54, 55], are inspired by the primate retina and operate asynchronously at the pixel level, providing ultra-high temporal resolution, wide dynamic range, low latency, and reduced energy consumption. Unlike event cameras, which follow a differential sampling model and capture only luminance changes in logarithmic space, spike cameras employ an integral sampling model, accumulating photons until a threshold is reached and then firing spikes. Although event cameras tend to generate sparser output, they might lose scene textures, especially in static regions due to their change-based design. In contrast, spike cameras can capture richer spatial details, making them well suited for tasks such as high-speed imaging [23, 26, 56, 57], 3D reconstruction [58], motion deblurring [43, 59], optical flow [24, 25], object detection [60], occlusion removal [61] and monocular depth estimation [21].
Frame-Based Stereo Depth Estimation. Deep learning has significantly advanced stereo matching, leading to various methods with enhanced accuracy and generalization. The early methods [62, 63, 64, 65] constructed cost volumes for each disparity and applied 3D CNNs for refinement. In contrast, recurrent refinement-based models [31, 32, 33, 34, 36] inspired by RAFT-Stereo [29, 49] iteratively update the disparity without building a full 4D volume. It restricts 2D flow to the 1D disparity dimension and uses multi-level convolutional Gated Recurrent Unit (ConvGRU) for broader receptive fields, demonstrating strong generalization. Successor models such as CREStereo [30], IGEV-Stereo [33] and Selective-Stereo [36] improve in-domain performance, but RAFT-Stereo retains advantages in zero-shot settings [29]. Recent hybrid methods [66] combine cost filtering and iterative refinement for better trade-offs. Meanwhile, transformer-based solutions [37] utilize cross-attention or global self-attention. As deep stereo matching evolves, fully leveraging vision foundation models remains an open opportunity for real-world applications.
Event-Based Stereo Depth Estimation. Event cameras capture pixel-level brightness changes asynchronously, offering significant advantages over conventional frame-based cameras. Recently, event-based stereo depth estimation has advanced rapidly, with notable methods exploiting camera velocity [45] or deriving depth without explicit event matching [67]. Deep learning approaches propose novel sequence embedding [68] and fuse frame and event data [69, 70] to improve depth estimates in challenging scenarios. Moreover, some works [71] leverage standard models from the image domain, exploiting the inherent connection between frame and event data to improve stereo depth estimation performance.
@@ 3 Methods
This section describes the overall architecture of our proposed model (Fig. 2). Figs. 2a, b illustrate the biologically inspired framework design, where stereo signals from both cameras are processed through a context network analogous to the visual pathway. Correlation features are computed and temporally refined via recurrent spiking neural network (RSNN) blocks, mimicking depth perception mechanisms in the brain. Fig. 2c shows the overall network architecture and consists of four main components: a multi-scale spike feature extractor, a correlation volume construction module, a biologically inspired RSNN-based update operator, and a final upsampling module for high resolution disparity refinement.
@@
The context network shares a similar backbone with the feature network. It consists of residual blocks and additional downsampling layers to produce multi-scale context features $f_{c,i}$ at 1/4, 1/8, and 1/16 of the original resolution, each with 128 channels. These features are used to initialize and inject context into the RSNNs at each recurrent iteration.
@@
Spatial Upsampling. Based on the hidden state of the final layer, the network predicts both residual disparity through two convolutional layers and then updates the current disparity: $d_{t}=d_{t-1}+\Delta d_{t}$. Finally, the disparity map at 1/4 resolution is upsampled to the full resolution using a convex combination strategy.
Loss Function. We supervise the network for a composite loss function composed of three terms: the main loss, the firing rate regularization, and the voltage regularization. The overall loss is defined as:
@@
$=\sum_{t=1}^{T}\eta^{T-t}\|d_{\text{gt}}-d_{t}\|_{1}+\lambda_{f}\sum_{i=1}^{N}(r_{i}-r_{0})^{2}+\lambda_{v}\sum_{i=1}^{N}\sum_{t=1}^{T}v_{i}(t)^{2},\quad\eta=0.9.$
The first term $L_{\text{stereo}}$ is the main loss, which is the $L_{1}$ norm distance between all predicted disparities $\left\{{d_{t}}\right\}^{T}_{t=1}$ and the ground truth disparity $d_{\text{gt}}$ with increasing weights. To constrain the firing rate of neurons and promote temporal sparsity, the second term $L_{\text{rate\_reg}}$ is firing rate regularization, $r_{i}$ is the average firing rate of the $i$-th neuron during time $T$, and $r_{0}$ is the target firing rate. Furthermore, the third term $L_{\text{v\_reg}}$ serves as the voltage regularization term, where $N$ represents the total number of neurons. Both regularization terms encourage sparsity and benefit the performance.
@@ 4.1 Datasets
Synthetic Dataset. To support supervised training for spike-based depth estimation, we construct a high-quality synthetic dataset using Blender [72]. The dataset includes 150 training scenes and 40 testing scenes, each containing 256 pairs of RGB images and the corresponding ground-truth depth maps. The dataset construction process [73, 74] and additional details are provided in the Supplementary Materials.
To generate dense spike streams with high temporal resolution in dynamic environments, we apply a video interpolation method [75] to synthesize 50 intermediate RGB frames between each pair of adjacent frames. These high-frequency image sequences are then converted into spike streams using a biologically inspired spike generation mechanism that models the analog behavior of real spiking camera circuits. Noise modeling is also integrated to further enhance realism. Our dataset provides synchronized high-resolution spike streams and pixel-wise depth ground truth, offering a comprehensive and reliable benchmark for training and testing.
Real Dataset. We collect a real-world spike stereo dataset using two spike cameras and one depth camera (Kinect) in diverse environments to evaluate the generalization and practical performance of the model. The spike cameras have a resolution of $400\times 250$ with a temporal resolution of 20,000 Hz, while the depth camera provides RGB-D data at $1280\times 720$ resolution and 30 FPS. The captured scenes involve various objects of the real world recorded in indoor settings. The dataset includes over 2000 stereo spike stream pairs (more than 200,000 frames) and depth maps, and is divided into training and test sets. Further details about our real dataset and the spike stereo binocular system are provided in the Supplementary Materials.
@@ 4.2 Implementation Details
We implement SpikeStereoNet using PyTorch and perform all experiments on NVIDIA RTX 4090 GPUs. The model is trained using the AdamW optimizer with gradient clipping in the range of $[-1,1]$. A one-cycle learning rate schedule is adopted, with an initial learning rate set to $2\times 10^{-4}$. During training, we apply 16 update iterations for each sample. We randomly cropped the spike stream to $224\times 384$ spatially during the training procedure and set the batch size as 8 for 300k steps. We use random horizontal and vertical flips as data augmentation to balance the motion in the training dataset.
@@ 4.3 Experimental Results
Our method is compared with various categories of stereo depth estimation approaches. First, event-based stereo networks are compared, such as ZEST [47]. Next, state-of-the-art frame-based stereo networks are evaluated, including Selective-Stereo [36], RAFT-Stereo [29], and GMStereo [37], etc. These methods cover different strategies, such as iterative refinement or transformer-based global matching. All networks are retrained using the same settings and evaluated under identical conditions to ensure fair comparison.
Experiments on the Synthetic Dataset. We use the average end-point error (AvgErr) and the bad pixel ratio with different pixel thresholds to evaluate the quality of stereo depth estimation. The results of the quantitative comparisons are presented in Table 1. As shown in the table, we achieve the best performance among the listed methods for almost all metrics in the synthetic dataset. Fig. 4 illustrates the scene, the spike stream, the ground truth of the depth, and the predicted depth maps of different methods. Our method significantly outperforms competing approaches, producing more accurate and sharper depth results. The sharpness and density of the predicted depth maps demonstrate the effectiveness of our approach in the spike stream input, especially in complex and cluttered scenes.
In addition to accuracy metrics, Table 1 also reports the params and floating-point operations (FLOPs) for each method. SpikeStereoNet achieves a favorable balance between performance and efficiency, ranking second in FLOPs while outperforming most models in parameter number. This demonstrates that the additional computational cost, incurred to better capture the spatiotemporal characteristics and asynchrony of spike data is justified. Compared to other methods that handle complex temporal structures, such as Transformer-based approaches, SpikeStereoNet remains highly competitive in terms of efficiency. Additional comparisons and analyses are provided in Supplementary Materials.
Table 1: Quantitative results on the testing set of synthetic dataset, and all methods use spike streams as input. Best results for each evaluation metric are bolded, second best are underlined. Lower is better for all metrics.
Method bad 1.0 $\downarrow$ bad 2.0 $\downarrow$ bad 3.0 $\downarrow$ AvgErr $\downarrow$ Params FLOPs
[…]
CREStereo [30] 16.12 10.54 4.40 0.81 5.43 863
RAFT-Stereo [29] 9.67 4.64 2.76 0.48 11.41 798
GMStereo [37] 17.61 11.05 4.39 0.83 7.40 160
IGEV-Stereo [33] 12.50 6.50 4.27 0.76 12.77 614
DLNR [31] 10.15 4.67 2.81 0.55 57.83 1580
ZEST [47] 11.10 4.94 3.50 0.62 340.75 989
MoCha-Stereo [39] 11.93 5.67 4.39 0.73 20.96 935
Selective-Stereo [36] 9.24 4.57 2.66 0.45 13.32 957
MonSter [40] 9.23 4.64 2.72 0.46 388.69 1567
SpikeStereoNet (Ours) 9.22 4.52 2.68 0.44 12.15 537
@@
Experiments on the Real Dataset. To bridge the domain gap between synthetic and real data, we apply a domain adaptation strategy to fine-tune our model, which is initially trained on synthetic datasets. The visualization results are shown in Fig. 5. Compared with other methods, our approach produces more accurate depth maps with sharper object boundaries and finer structural details. It performs especially well in challenging regions, such as textureless surfaces and high-illumination areas, benefiting from the integration of spike domain characteristics. These results highlight the effectiveness of our domain adaptation strategy and the robustness of our model in real-world stereo depth estimation from spike streams. More quantitative comparison experiments and analyses are accessible in the Supplementary Materials.
Figure 5: Visual results of our method and competing approaches on the real dataset. The "Scene" refers to the gamma-transformed temporal average of spike streams. "Kinect" represents the raw depths captured by the depth camera.
@@
Explanation of Iterative Refinement Structures. We analyze the dynamic behavior of our hybrid neural network to assess its temporal characteristics. As shown in Fig. 7, the left panel illustrates that hidden state differences decrease over time, indicating convergence. The middle panel displays the eigenvalue spectrum of the Jacobian matrix at the final time step, with all eigenvalues lying within the unit circle, which confirms the stability of the system. The right panel presents a PCA of hidden states across input batches, showing increasing dispersion over time, which reflects the model’s ability to encode diverse temporal patterns. These results highlight the temporal stability, robustness, and expressive power of the network.
Beyond validating our specific RSNN design, this dynamics analysis sheds light on why iterative refinement, the core mechanism in architectures like RAFT-Stereo [29], is so effective. Our results explicitly show how a well-behaved iterative updater can maintain stability while progressively integrating information over time steps to refine the estimate towards the correct solution. This provides a dynamical grounding for the empirical success of iterative methods.

### v2

@@ 1 Introduction
Depth perception is fundamental for navigating and interacting with the 3D world (Tosi et al., 2025), driving applications from robotic manipulation (Ma et al., 2024; Li et al., 2024; Wen et al., 2025b; Gao et al., 2024) to autonomous navigation (Wei et al., 2024; Duba et al., 2024; Kalenberg et al., 2024; Nahavandi et al., 2022; Wu et al., 2023; Xu et al., 2023c; Wang et al., 2024b). Traditional stereo vision estimates depth from calibrated image pairs captured by frame-based cameras; however, it suffers from motion blur and latency in dynamic scenes. Biological vision systems efficiently process visual inputs using sparse, asynchronous spikes, achieving remarkable speed and energy efficiency (Kundu et al., 2021; Yang et al., 2023; Datta et al., 2021; Göltz et al., 2021; Wang et al., 2025; Wolf & Lappe, 2021; Kucik & Meoni, 2021). Neuromorphic spike cameras (Zhao et al., 2024; Zhang et al., 2022a; Zhao et al., 2021a; b; Hu et al., 2022; Zhao et al., 2022; Chen et al., 2024a) implement this principle in hardware, delivering ultra-high temporal resolution (up to 40,000 Hz) and capturing rich luminance information through asynchronous binary spike streams. This makes them suitable for robust perception tasks, such as stereo depth estimation, particularly in highly dynamic scenarios where conventional methods are ineffective.
@@
Despite their potential, spike cameras present distinct challenges for stereo depth estimation. Their asynchronous, binary, high‑throughput streams conflict with frame‑based algorithms that expect synchronous, intensity‑valued image pairs (Tankovich et al., 2021; Lipson et al., 2021; Li et al., 2022; Zhao et al., 2023; Rao et al., 2023; Xu et al., 2023a; Zeng et al., 2023; Guan et al., 2024; Wang et al., 2024a; Xu et al., 2023b; Weinzaepfel et al., 2023; Chen et al., 2024b; Cheng et al., 2025; Jiang et al., 2025; Wen et al., 2025a), and any conversions to frames introduce temporal quantization errors, motion blur, or significant computational overhead (Zhao et al., 2021b; 2022; Chen et al., 2023), diminishing the sensor’s intrinsic advantages. Although event‑based stereo methods have been developed for DVS cameras (e.g. (Zhou et al., 2021; Zhang et al., 2022b; Cho et al., 2021; Lou et al., 2024; Rançon et al., 2022)), they rely on temporal contrast, while spike cameras emit integrated intensity streams that require tailored processing techniques. Critically, the field lacks specialized algorithms and benchmarks designed for stereo depth estimation directly from raw spike streams, hindering progress in this promising field.
@@
To bridge this gap, we introduce SpikeStereoNet (Fig. 1), an end-to-end, biologically inspired framework for generating stereo depth directly from raw spike streams. SpikeStereoNet integrates a recurrent spiking neural network (RSNN) into an iterative refinement loop (Teed & Deng, 2020) and models neuronal interactions (Kandel et al., 2021) to capture spatiotemporal dynamics inherent in spike data. The analysis of neuronal dynamics confirms the temporal stability, convergence, and expressive feature separation of the RSNN components. We introduce two novel benchmark datasets: a large-scale synthetic dataset with diverse scenes and ground-truth depth, and a real-world dataset containing synchronized stereo spike streams and corresponding depth sensor measurements. Experimental results demonstrate that SpikeStereoNet generalizes effectively, maintaining robust performance even when trained with substantially limited data. Our main contributions are as follows:
@@ 2 Related Work
Spike Cameras and its Applications. Neuromorphic cameras, including event cameras (Moeys et al., 2017; Posch et al., 2010; Huang et al., 2017) and spike cameras (Huang et al., 2023), are inspired by the primate retina and operate asynchronously at the pixel level, providing ultra-high temporal resolution, wide dynamic range, low latency, and reduced energy consumption. Unlike event cameras, which follow a differential sampling model and capture only luminance changes in logarithmic space, spike cameras employ an integral sampling model, accumulating photons until a threshold is reached and then firing spikes. Although event cameras tend to generate sparser output, they might lose scene textures, especially in static regions due to their change-based design. In contrast, spike cameras can capture richer spatial details, making them well suited for tasks such as high-speed imaging (Zhu et al., 2020), 3D reconstruction (Zhang et al., 2024b), motion deblurring (Zhang et al., 2024c), optical flow (Zhao et al., 2022), object detection (Zheng et al., 2022), occlusion removal (Zhang et al., 2023b) and monocular depth estimation (Zhang et al., 2022a).
Frame-Based Stereo Depth Estimation. Deep learning has significantly advanced stereo matching, leading to various methods with enhanced accuracy and generalization. The early methods (Fan et al., 2021; Yin et al., 2019) constructed cost volumes for each disparity and applied 3D CNNs for refinement. In contrast, recurrent refinement-based models (Rao et al., 2023; Tian et al., 2023) inspired by RAFT-Stereo (Lipson et al., 2021; Teed & Deng, 2020) iteratively update the disparity without building a full 4D volume. It restricts 2D flow to the 1D disparity dimension and uses multi-level convolutional Gated Recurrent Unit (ConvGRU) for broader receptive fields, demonstrating strong generalization. Successor models such as DLNR (Zhao et al., 2023), IGEV-Stereo (Xu et al., 2023a) and Selective-Stereo (Wang et al., 2024a) improve in-domain performance, but RAFT-Stereo retains advantages in zero-shot settings (Lipson et al., 2021). Recent hybrid methods (Hamid et al., 2021) combine cost filtering and iterative refinement for better trade-offs. Meanwhile, transformer-based solutions (Xu et al., 2023b) utilize cross-attention or global self-attention.
Event-Based Stereo Depth Estimation. Event cameras capture pixel-level brightness changes asynchronously, offering significant advantages over conventional frame-based cameras. Recently, event-based stereo depth estimation has advanced rapidly (Gallego et al., 2020; Tosi et al., 2025), with notable methods exploiting camera velocity (Zhang et al., 2022b) or deriving depth without explicit event matching (Zhou et al., 2018). Deep learning approaches propose novel sequence embedding (Tulyakov et al., 2019) and fuse frame and event data (Nam et al., 2022) to improve depth estimates in challenging scenarios. Moreover, some works (Cho et al., 2023) leverage standard models from the image domain, exploiting the inherent connection between frame and event data to improve stereo depth estimation performance.
@@ 3 Methods
This section describes the overall architecture of our proposed model (Fig. 2). Figs. 2a, b illustrate the design of the biologically inspired framework. Fig. 2c shows the overall network architecture, which consists of the following four main components.
@@
The context network shares a similar backbone with feature networks. It consists of residual blocks and additional downsampling layers to produce multi-scale context features of the original resolution. These features are used to initialize and inject contexts into RSNNs at each recurrent iteration.
@@
Spatial Upsampling. Based on the hidden state of final layer, network predicts both residual disparity through two convolutional layers and then updates current disparity: $d_{t}=d_{t-1}+\Delta d_{t}$. Finally, the disparity at 1/4 resolution is upsampled to full resolution using a convex combination strategy.
Loss Function. We supervise the network for a composite loss function composed of the three terms: the main loss, the firing rate and the voltage regularization. The overall loss is defined as:
@@
[…]
The first term $L_{\text{stereo}}$ is the main loss, which is the $L_{1}$-norm distance between all predicted disparities $\left\{{d_{t}}\right\}^{T}_{t=1}$ and the ground truth disparity $d_{\text{gt}}$ with increasing weights. To constrain the firing rate of neurons and promote temporal sparsity, the second term $L_{\text{rate\_reg}}$ is firing rate regularization, $r_{i}$ is the average firing rate of the $i$-th neuron during time $T$, and $r_{0}$ is the target firing rate. Furthermore, the third term $L_{\text{v\_reg}}$ serves as the voltage regularization term, where $N$ represents the total number of neurons. Both regularization terms encourage sparsity and benefit the performance.
@@ 4.1 Datasets
Synthetic Dataset. To support supervised training for spike-based depth estimation, we construct a high-quality synthetic dataset using Blender (Community, 2018). The dataset includes 150 training scenes and 40 testing scenes, each containing 256 pairs of RGB images and corresponding ground-truth depth maps. The dataset build process and details are in the Appendix A.3.
To generate dense spike streams with high temporal resolution in dynamic environments, we apply a video interpolation method (Zhang et al., 2023a) to synthesize 50 intermediate RGB frames between each pair of adjacent frames. These high-frequency image sequences are then converted into spike streams using a biologically inspired spike generation mechanism that models the analog behavior of real spiking camera circuits. Noise modeling is also integrated to further enhance realism. Our dataset provides synchronized high-resolution spike streams and pixel-wise depth ground truth, offering a comprehensive and reliable benchmark for training and testing.
Real Dataset. We collect a real-world spike stereo dataset using two spike cameras and one depth camera (Kinect) in diverse environments to evaluate the generalization and practical performance of the model. The spike cameras have a resolution of $400\times 250$ with a temporal resolution of 20,000 Hz, while the depth camera provides RGB-D data at $1280\times 720$ resolution and 30 FPS. The captured scenes involve various objects of the real world recorded in indoor settings. The dataset includes over 3,000 stereo spike stream pairs (more than 300,000 frames) and depth maps and is divided into training and test sets. More details about our real dataset and system are provided in Appendix A.4.
@@ 4.2 Implementation Details
We implemented SpikeStereoNet using PyTorch and performed all experiments on NVIDIA RTX 4090 GPUs. The model is trained using the AdamW optimizer with gradient clipping in the range of [1,1][-1,1]. A one-cycle learning rate schedule is adopted, with an initial learning rate set to $2\times 10^{-4}$. During training, we apply 16 update iterations for each sample and set the batch size as 8 for 300k steps. We use random horizontal and vertical flips as data augmentation in the training dataset.
@@ 4.3 Experimental Results
Our method is compared with various categories of stereo depth estimation approaches. First, event-based stereo networks are compared, such as ZEST (Lou et al., 2024). Next, state-of-the-art frame-based stereo networks are evaluated, including Selective-Stereo (Wang et al., 2024a), RAFT-Stereo (Lipson et al., 2021), and GMStereo (Xu et al., 2023b), etc. These methods cover different strategies, such as iterative refinement or transformer-based global matching. All networks are re-trained using the same settings and evaluated under identical conditions to ensure fair comparison.
Table 1: Quantitative results on the test set of synthetic dataset, and all methods use spike streams as input. Best results for each evaluation metric are bolded, second best are underlined. The $\downarrow$” indicates that the lower the metrics, the better.
Method bad 1.0 bad 2.0 bad 3.0 AvgErr Params FLOPs
$(\%)\downarrow$ $(\%)\downarrow$ $(\%)\downarrow$ (px) $\downarrow$ (M) (B)
CREStereo (Li et al., 2022) 16.12 10.54 4.40 0.81 5.43 863
Stereospike (Rançon et al., 2022) 14.10 9.82 5.35 1.10 5.10 569
RAFT-Stereo (Lipson et al., 2021) 9.67 4.64 2.76 0.48 11.41 798
GMStereo (Xu et al., 2023b) 17.61 11.05 4.39 0.83 7.40 160
IGEV-Stereo (Xu et al., 2023a) 12.50 6.50 4.27 0.76 12.77 614
DLNR (Zhao et al., 2023) 10.15 4.67 2.81 0.55 57.83 1580
ZEST (Lou et al., 2024) 11.10 4.94 3.50 0.62 340.75 989
MoCha (Chen et al., 2024b) 11.93 5.67 4.39 0.73 20.96 935
Selective (Wang et al., 2024a) 9.24 4.57 2.66 0.45 13.32 957
MonSter (Cheng et al., 2025) 9.23 4.64 2.72 0.46 388.69 1567
SpikeStereoNet (Ours) 8.41 4.13 2.38 0.42 12.15 473
@@
Experiments on the Synthetic Dataset. We use the average end-point error (AvgErr) and the bad pixel ratio with different pixel thresholds to evaluate the quality of stereo depth estimation. The results of the quantitative comparisons are presented in Table 1. As shown in the table, we achieve the best performance among the listed methods for almost all metrics in the synthetic dataset. Fig. 4 illustrates the scene, the spike stream, the ground truth of the depth, and the predicted depth maps of different methods. Our method significantly outperforms competing approaches, producing more accurate and sharper depth results. The sharpness and density of predicted depth maps demonstrate the effectiveness of our approach in spike stream inputs, especially in complex and cluttered scenes.
In addition to accuracy metrics, Table 1 also reports the params and floating-point operations (FLOPs) for each method. SpikeStereoNet achieves a favorable balance between performance and efficiency, ranking second in FLOPs while outperforming most models in parameter number. This demonstrates that the additional computational cost, incurred to better capture the spatiotemporal characteristics and asynchrony of spike data is justified. Compared to other methods that handle complex temporal structures, such as Transformer-based approaches, SpikeStereoNet remains highly competitive in terms of efficiency. Additional comparison results are provided in Appendix.
Experiments on the Real Dataset. To bridge the domain gap between synthetic and real data, we apply a domain adaptation strategy to fine-tune our model, which is initially trained on synthetic datasets. The visualization results are shown in Fig. 5. Compared with other methods, our approach produces more accurate depth maps with sharper object boundaries and finer structural details. It performs especially well in challenging regions, such as textureless surfaces and high-illumination areas, benefiting from the integration of spike domain characteristics. These results highlight the effectiveness of our domain adaptation strategy and the robustness of our model in real-world stereo depth estimation from spike streams. More quantitative results can be found in the Appendix A.5.
Figure 5: Visual results of our method and competing approaches on the real dataset. The “Scene” refers to the gamma-transformed temporal average of spike streams. “Kinect” represents the raw depths captured by the depth camera.
@@
Explanation of Iterative Refinement Structures. We analyze the dynamic behavior of our hybrid neural network to assess its temporal characteristics. As shown in Fig. 7, the left panel illustrates that hidden state differences decrease over time, indicating convergence. The middle panel displays the eigenvalue spectrum of the Jacobian matrix at the final time step, with all eigenvalues located within the unit circle, which confirms the stability of the system. The right panel presents a PCA of hidden states across input batches, showing increasing dispersion over time, which reflects the model’s ability to encode diverse temporal patterns. These results highlight temporal stability, robustness, and expressive power of the network. Further theoretical proof can be found in Appendix A.5.
Beyond validating our specific RSNN design, this dynamics analysis sheds light on why iterative refinement, the core mechanism in architectures like RAFT-Stereo (Lipson et al., 2021), is so effective. Our results explicitly show how a well-behaved iterative updater can maintain stability while progressively integrating information over time steps to refine the estimate towards the correct solution. This provides a dynamical grounding for the empirical success of iterative methods.